Haas Company manufactures and sells one product. The following information pertains to each of the company’s first three years of operations:

Variable costs per unit:
Manufacturing:
Direct materials $ 21
Direct labor $ 13
Variable manufacturing overhead $ 4
Variable selling and administrative $ 2
Fixed costs per year:
Fixed manufacturing overhead $ 330,000
Fixed selling and administrative expenses $ 150,000

During its first year of operations, Haas produced 40,000 units and sold 40,000 units. During its second year of operations, it produced 55,000 units and sold 30,000 units. In its third year, Haas produced 20,000 units and sold 45,000 units. The selling price of the company’s product is $52 per unit.

Required:

1. Compute the company’s break-even point in unit sales.

2. Assume the company uses variable costing:

a. Compute the unit product cost for Year 1, Year 2, and Year 3.

b. Prepare an income statement for Year 1, Year 2, and Year 3.

3. Assume the company uses absorption costing:

a. Compute the unit product cost for Year 1, Year 2, and Year 3.

b. Prepare an income statement for Year 1, Year 2, and Year 3.
Answer #1

1) Break even point(BEP)(in unit sales) = Total fixed costs / Contribution per unit *(all amounts in $)

fixed costs = 330000+150000 =480000

contribution per unit = selling price - (direct materials+direct labour+ variable expenses)

= 52-(21+13+4+2) =12

hence, BEP = 480000/12

=40000 units.

2)

a) Unit product cost

for year 1 - (direct materials + direct labour + variable expenses)

= 21+13+4+2

=40

The unit product cost will be same for all the three years i.e., 40 since they are fixed.

b) income statement

for year 1   

sales 52

(-) variable costs 40

contribution 12

total contribution 480000(40000*12)

(-) fixed costs 480000(330000+150000)

profit 0

for year 2   

sales 52

(-) variable costs 40

contribution 12

total contribution 360000(30000*12)

(-) fixed costs 480000(330000+150000)

loss 120000

for year 3   

sales 52

(-) variable costs 40

contribution 12

total contribution 540000(45000*12)

(-) fixed costs 480000(330000+150000)

profit 60000

3)

a) Unit product cost

for year 1 - (direct materials + direct labour + variable expenses + per unit fixed costs)

= 21+13+4+2+(330000/40000)+(150000/40000) [fixed manufacturing OH is related to production and fixed selling and administrative expenses is related to sales]

=52

for year 2 - (direct materials + direct labour + variable expenses + per unit fixed costs)

= 21+13+4+2+(330000/55000)+(150000/30000) [fixed manufacturing OH is related to production and fixed selling and administrative expenses is related to sales]

=51

for year 3 - (direct materials + direct labour + variable expenses + per unit fixed costs)

= 21+13+4+2+(330000/20000)+(150000/45000) [fixed manufacturing OH is related to production and fixed selling and administrative expenses is related to sales]

=59.83  

b) income statement

for year 1   

sales 52

product cost 52

profit per unit 0

hence total profit is 40000*0=0

for year 2   

sales 52

product cost 51

profit per unit 1

hence total profit is 30000*1=30000

for year 3   

sales 52

product cost 59.83

loss per unit 7.83

hence total loss is 45000*7.83=352350